OpenStudy (anonymous):

integral of x^3/(sqrt(16-x^2))

OpenStudy (raden):

let u = 16 - x^2 ---------> x^2 = (16-u) du = -2x dx or -1/2 du = x dx see ur integral becomes int (-1/2 (16-u))/sqrt(u) du = int -8/sqrt(u) du + int (1/2 u)/sqrt(u) du = int -8(u)^-1/2 du + int 1/2 u^1/2 du now, integrate this

OpenStudy (raden):

use the basic formula :